OpenStudy (anonymous):

At 9:00 on Saturday morning, two bicyclists heading in opposite directions pass each other on a bicycle path.The bicyclist heading north is riding 4 km/hour faster than the bicyclist heading south. At 10:30, they are 39 km apart. Find the two bicyclists’ rates. (A)northbound bicyclist = 15 km/h; southbound bicyclist = 11 km/h (B)northbound bicyclist = 17 km/h; southbound bicyclist = 11 km/h (C)northbound bicyclist = 15 km/h; southbound bicyclist = 10 km/h (D)northbound bicyclist = 16 km/h; southbound bicyclist = 12 km/h

OpenStudy (e.cociuba):

D.) Let me explain...

OpenStudy (anonymous):

when i tried the problem i got C

OpenStudy (e.cociuba):

So 1.5hours*(2x+8 km/hour)=48 km 2x+8=48/1.5 2x+8=32 2x=24 x=12 This is wat I got...

OpenStudy (e.cociuba):

Is tht rite @Mertsj ?

OpenStudy (anonymous):

I was going to answer, but @e.cociuba with here you are in good hands

OpenStudy (e.cociuba):

Hahah thanks @Luis_Rivera

OpenStudy (mertsj):

Should be 2r + 4

OpenStudy (anonymous):

thank u soo much

OpenStudy (e.cociuba):

So I did It wrong...

OpenStudy (mertsj):